# FSAG: Enhancing Human-to-Dexterous-Hand Finger-Specific Affordance Grounding via Diffusion Models

### v1

@@ IV-C Functional Affordance Grasping with High-DOF Hands
Our evaluation targets on-table, affordance-driven dexterous grasping with finger-specific contacts and human-consistent grasp sites. Existing simulation-trained policies [36] optimize for success in simplified settings and cannot select grasp sites following human conventions (e.g., grasping a hammer by its head). To assess alignment with human grasp choices within our setting, we include two point-cloud-based imitation learning policies (Diffusion Policy 3D [42] and ACT-3D [11]) trained from our teleoperated demonstrations(30 trajectories). We also compare to CMKA representing a recent affordance-based dexterous grasping approach. We evaluate CMKA strictly under its original, unmodified pipeline: we directly use the three keypoints predicted by CMKA, lift them to 3D, assign them to the functional finger, little finger, and wrist, and then execute the method’s fixed grasp execution strategy.
@@
As shown in Tab. II, the conventional imitation-learning baselines tend to reproduce trajectories close to demonstrations: as object morphology varies, they do not proactively adapt their strategy, leading to significantly degraded performance both on seen objects with poses absent from the training data and on unseen objects. CMKA’s performance is limited by the layered segmentation output of SAM; as discussed in Section II-B, for objects that lack salient part delineation such as tomatoes and bananas, it often fails to predict correct grasp locations. Moreover, because it executes a preprogrammed grasp approaching motion, the fingers collide with the table surface when attempting to grasp low-profile objects such as bananas and screwdrivers, thus preventing successful task completion. By contrast, our method achieves higher success on common objects and remains robust on previously unseen instances within the same categories. In Fig. 6, per-finger affordance inference places contacts on grasp-functional geometry, while robust stereo depth estimation provides stable surface normals for accurate on-table execution. The geometry-aligned Approach Vector co-optimizes wrist pose and finger trajectories relative to the tabletop object, producing table-skimming entries that reliably avoid finger–table collisions, particularly beneficial for low-profile objects.
@@
TABLE II: Performance comparison across different methods on common objects and tools manipulation tasks. Each cell reports N=20N{=}20 trials per object (S = seen, U = unseen).
@@
Ours 100/85 85 90 60 70
IV-D Cross-Embodiment Evaluation
@@
As shown in the right column of Fig. 1, we maintain the same perception-to-optimization pipeline across dexterous-hand embodiments, performing finger-specific affordance inference fused with depth, followed by an approach–closure–hold planner. We transfer the pipeline from DexHand021 (12+512{+}5 DOFs) to Linker Hand L20 (16+516{+}5 DOFs) via kinematic retargeting of fingertip targets; perception and affordance inference remain unchanged. As shown in Tab. III, grasp success remains statistically unchanged across embodiments. This invariance indicates that the proposed finger-specific affordance representation and geometry-aligned waypoint construction decouple grasp semantics from hand morphology, enabling deployment without hand-specific retraining.
TABLE III: Grasp success rates across different hand embodiments.
Common Objects (%) Tools (%)
Model Bottle (S/U) Banana Tomato Screwdriver Drill
DexHand021 100/85 85 90 60 70
Linker Hand L20 100/80 90 100 70 90
@@ V CONCLUSION AND FUTURE WORK
We leverage pretrained generative models’ understanding of hand–object interactions to extract both high-level object semantics and fine-grained visual cues, and fuse these with a small set of human demonstrations providing fingertip annotations. This enables learning finger-specific affordances that unify perception and control: our representation captures not only where to act but also how to act. Built on this, our FSAG-based motion planning maps dexterous-hand postures onto finger-specific contacts, yielding stable functional grasps and generalization across multiple dexterous-hand embodiments. Empirically, our approach substantially outperforms contemporary methods on both affordance localization accuracy and functional grasping success rate.
As illustrated in Fig. 7, occasional failures on slender objects arise primarily from the execution: a simple, fixed-offset open-loop closure can induce rotation or slip even when contact targets are valid. Next, we will close the loop around FSAF contacts via learned residual-impedance/operational-space control with contact-force estimation and slip detection, yielding more robust and dynamic manipulation.

### v2

@@ IV-C Functional Affordance Grasping with High-DOF Hands
Our evaluation targets on-table, affordance-driven dexterous grasping with finger-specific contacts and human-consistent grasp sites. Existing simulation-trained policies [22] optimize for success in simplified settings and cannot select grasp sites following human conventions (e.g., grasping a hammer by its head). To assess alignment with human grasp choices within our setting, we include two point-cloud-based imitation learning policies (Diffusion Policy 3D [26] and ACT-3D [7]) trained from our teleoperated demonstrations (30 trajectories). We also compare to CMKA, representing a recent affordance-based dexterous grasping approach. We evaluate CMKA strictly under its original, unmodified pipeline: we directly use the three keypoints predicted by CMKA, lift them to 3D, assign them to the functional finger, little finger, and wrist, and then execute the method’s fixed grasp execution strategy. We additionally compare with GrainGrasp [27], a representative fine-grained contact-guided grasp synthesis method, by adapting its contact-guided refinement stage to our robotic hand embodiment and evaluating it under the same perception and execution pipeline for real-robot success-rate comparison.
Fig. 6 presents qualitative grasp predictions on the Linker Hand L20 for GrainGrasp, GraspQP, and our method. Our approach consistently produces human-intuitive grasp poses: guided by finger-specific affordances, the fingertips align tightly with the predicted graspable regions, yielding contact layouts that match the functional geometry of each object. In contrast, GraspQP with random initialization performs markedly worse, often converging to non-human-like poses (e.g., on the screwdriver), missing the intended grasp regions, and occasionally producing degenerate configurations with overlapping fingers. GraspQP* (GraspQP with the same region-conditioned initialization as our method) can sometimes achieve similarly reasonable grasps on bottle, tomato, and screwdriver, indicating that a strong warm start already alleviates some local-minima failures; however, it still fails on the banana and converges to visually implausible configurations on the drill, suggesting that initialization alone is insufficient to enforce affordance-consistent, finger-wise contact placement. It’s also noteworthy that GrainGrasp fails to generate meaningful grasps across these objects, as it is tailored to MANO-scale human hands and its contact representation and kinematic assumptions do not transfer well to high-DOF robotic hands.
@@
As shown in Tab. II, the conventional imitation-learning baselines (ACT-3D AND Diffusion Policy 3D) tend to reproduce trajectories close to demonstrations: as object morphology varies, they do not proactively adapt their strategy, leading to significantly degraded performance both on seen objects with poses absent from the training data and on unseen objects. GrainGrasp, while leveraging contact cues extracted from human-hand demonstrations, is hindered by the lack of cross-demonstration consistency in the resulting contact patterns, which makes contact-conditioned pose synthesis brittle under distribution shift. Moreover, the method is evaluated only in simulation and its contact-to-grasp stage does not explicitly enforce real-robot grasp feasibility (e.g., force-closure and frictional stability under hardware and contact uncertainties); consequently, it attains near-zero success in our real-robot protocol (Tab. II), underscoring the practical difficulty of sim-to-real transfer for contact-guided grasp synthesis. Meanwhile, CMKA’s performance is limited by the layered segmentation output of SAM; as discussed in Section II-B, for objects that lack salient part delineation such as tomatoes and bananas, it often fails to predict correct grasp locations. Moreover, because it executes a preprogrammed grasp approaching motion, the fingers collide with the table surface when attempting to grasp low-profile objects such as bananas and screwdrivers, thus preventing successful task completion.
@@
TABLE II: Performance comparison across different methods on common objects and tools manipulation tasks. Each cell reports N=20N{=}20 trials per object (S = seen, U = unseen). We additionally report cross-embodiment transfer of our method on two dexterous hands.
@@
Ours (DexHand021) 100/85 85 90 60 70
Ours (Linker Hand L20) 100/80 90 100 70 90
Figure 8: Energy convergence over the full optimization. Energy vs. optimization step throughout the entire refinement process. The blue curve corresponds to GraspQP* (GraspQP with the same initialization as our method), while the red curve denotes our method.
Fig. 8 plots energy over the full refinement, including multiple resets. Although our objective augments the original GraspQP formulation with an additional affordance alignment term $E_{\mathrm{aff}}$, it consistently converges faster than GraspQP* after each reset and reaches a lower final energy overall. We attribute this improvement to the fact that $E_{\mathrm{aff}}$ provides a dense, region-consistent geometric signal that pulls finger contacts toward the predicted graspable surface regions, thereby shrinking the effective search space and steering the optimizer away from poor local minima. In practice, this contact-level guidance also indirectly benefits the force-closure term $E_{\mathrm{fc}}$: by promoting more plausible and better-situated contacts, the solver more readily finds feasible contact force distributions that balance external wrenches, resulting in higher-quality grasps at convergence.
Cross-Embodiment Evaluation.
Beyond the DexHand021 used in the preceding experiments, we further test the same pipeline on another dexterous hand embodiment (Linker Hand L20).We transfer the pipeline from DexHand021 (12+512{+}5 DOFs) to Linker Hand L20 (16+516{+}5 DOFs); perception and affordance inference remain unchanged. As shown in Tab. II, grasp success remains statistically unchanged across embodiments. This invariance indicates that the proposed finger-specific affordance representation and geometry-aligned waypoint construction decouple grasp semantics from hand morphology, enabling deployment without hand-specific retraining.
@@ V CONCLUSION AND FUTURE WORK
We propose Finger-Specific Affordance Grounding (FSAG), which learns per-finger contact affordances from a small set of human demonstrations. Our key idea is to repurpose a frozen text-to-image diffusion model as a semantic backbone and train a lightweight decoder to predict five finger-conditioned likelihood maps, bridging high-level interaction priors with fine-grained contact localization. We lift the predicted per-finger affordance maps to object-surface regions and couple them with physically grounded grasp refinement, jointly optimizing contact feasibility and finger–region consistency to produce stable, human-intuitive grasps. Empirically, our approach substantially outperforms contemporary methods on both affordance localization accuracy and human-intuitive grasping success rate.